OpenStudy (anonymous):

2sin^2(x)=2+cos(x)............i am so lost

OpenStudy (anonymous):

sub.s in sin^2(x) = 1-cos^2(x) I guess you don't really need the steps with 'u=cos(x)' in this problem...

OpenStudy (anonymous):

2(1-cos^2(x) ) = 2+cos(x) 2 -2cos^2(x) = 2+cos(x) -2cos^2(x) = cos(x) cos(x) = -1/2 converting it into a quadratic will come in handy down the road though, so keep it in mind....

OpenStudy (anonymous):

how did you get rid of -2cos^2(x)

OpenStudy (anonymous):

divide both sides by -2cos(x)